solve: y''(x^2 +1) = 2xy' y' = ist derivative y'' = second derivate!

solve:


y''(x^2 +1) = 2xy'


 


y' = ist derivative


y'' = second derivate!

Grade:12th Pass

2 Answers

Arun Kumar IIT Delhi
9 years ago
substitue y' =u
y''=u'
we get u'=2x/(x*x+1)u
ln(u)=ln(x^2+1)+C
=> u=c(x^2+1)
dy/dx=c(x^2+1)
y=c((x^3)/3+x)+c1
TEJKIRAT SINGH
9 years ago
substitue y' =uy''=u'we get u'=2x/(x*x+1)uln(u)=ln(x^2+1)+C=> u=c(x^2+1)dy/dx=c(x^2+1)y=c((x^3)/3+x)+c1